OpenStudy (cinnamonbun):

(1 pt) Vicki measured the midnight temperature at four locations in Alaska. The temperatures, in degrees Fahrenheit, were –7, –2, 1, and 4. What is the mean absolute deviation of the temperatures?

jimthompson5910 (jim_thompson5910):

First add up all the numbers. Then divide that result by 4. What do you get?

OpenStudy (cinnamonbun):

-1?

OpenStudy (cinnamonbun):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

that's the mean

jimthompson5910 (jim_thompson5910):

Set up a number line. Plot the mean -1 on the number line

jimthompson5910 (jim_thompson5910):

Focus on the first data value -7 and the mean -1 How far is -7 from -1 on the number line?

OpenStudy (cinnamonbun):

6?

jimthompson5910 (jim_thompson5910):

yep now focus on the data value -2 (second value) and the mean -1 how far is -2 from the mean -1 ?

OpenStudy (cinnamonbun):

1

jimthompson5910 (jim_thompson5910):

yep

jimthompson5910 (jim_thompson5910):

now focus on the data value 1 (third value) and the mean -1 how far is 1 from the mean -1 ?

OpenStudy (cinnamonbun):

2

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

how far is 4 (last data value) from the mean -1 ?

OpenStudy (cinnamonbun):

-4 from 1 is 5?

OpenStudy (cinnamonbun):

5

OpenStudy (cinnamonbun):

4 from -1 sorry

jimthompson5910 (jim_thompson5910):

So we have these distances 6 1 2 5 Add up those distances 6+1+2+5 = 14 Then divide that sum by 4 (because there are 4 values in the original data set) 14/4 = 7/2 = 3.5 So the mean absolute deviation (MAD) is 3.5 What does the MAD represent? It is the average distance each point is from the mean. It's also known as the mean deviation http://www.mathsisfun.com/data/mean-deviation.html

OpenStudy (cinnamonbun):

Thank you so so much!!!

jimthompson5910 (jim_thompson5910):

The larger the MAD is, the more spread out the data is. The smaller the MAD, the more clumped the data is.

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (cinnamonbun):

I see, ty!